# When Can Model-Free Reinforcement Learning be Enough for Thinking?

### v1

@@ Appendix A Proofs of Theoretical Results
The proof follows from the fact that, under the assumed policy initialization, thought actions raise the lower bound on finding a goal. First, note that the effective horizon has a tighter upper bound if we always fix the thought state to be $\tau_{1}$ rather than $\tau_{0}$. Thus, there is just the question of getting the thought state to be $\tau_{1}$. We lower bound this probability with $p_{c}$ so that, if the agent is in $(s,\tau_{0})$, it has at least a probability of getting to $\tau_{1}$ from which it has at least a probability $p_{1}$ of finding a goal. Thus, the lower bound on $\Pr(s_{T}\in\mathcal{S}_{\mathtt{goal}}|s,\tau,a,\pi)$ will be at least $p_{c}\cdot p_{1}$ for both $\tau_{0}$ and $\tau_{1}$ which reduces the effective horizon if $p_{c}\cdot p_{1}>p_{0}$.
Appendix B Example Prompts for LLM Experiment
@@
Appendix C Extended Empirical Description of Gridworld Experiments
@@
The objective of our experiment described in Section 6 was to create a setting where model-free RL (in this case REINFORCE) would lead to thinking as a strategy for reward maximization. A crucial piece of the set-up was to have a pre-trained model for which thought actions increase the agent’s probability of solving the task. Recall that solving the task requires the agent to navigate to the bottom right corner and then to reach the top left corner. Pre-training is designed so that the agent outputting special action ’A’ increases the probability of it then taking the actions that lead to the bottom right and similarly so that ’B’ increases the probability of moving to the top left. We validate the pre-training procedure by taking pre-trained models and forcing their first action to be ’A’ and their action upon reaching the bottom right for the first time to be ’B’. We find that the pre-training procedure leads to models where such prompting increases the probability of task success compared to simply rolling out the model. If we further prompt the model this way and also mask thought actions on every other step then the pre-trained model will solve the task on virtually every episode. This prompting and masking procedure confirmed that pre-training had primed the model for thinking to emerge as an effective strategy.

### v2

@@ Appendix A Proofs of Theoretical Results
The proof follows from the fact that, under the assumed policy initialization, thought actions raise the lower bound on finding a goal. First, note that the effective horizon has a tighter upper bound if we always fix the thought state to be $\tau_{1}$ rather than $\tau_{0}$. Thus, there is just the question of getting the thought state to be $\tau_{1}$. We lower bound this probability with $p_{c}$ so that, if the agent is in $(s,\tau_{0})$, it has at least a probability of getting to $\tau_{1}$ from which it has at least a probability $p_{1}$ of finding a goal. Thus, the lower bound on $\Pr(s_{T}\in\mathcal{S}_{\mathtt{goal}}|s,\tau,a,\pi)$ will be at least $p_{c}\cdot p_{1}$ for both $\tau_{0}$ and $\tau_{1}$ which reduces the effective horizon if $p_{c}\cdot p_{1}>p_{0}$. ∎
Appendix B Extended Theoretical Analysis for Finite-Horizon Thought MDPs
In the main paper text we formalized thought MDPs under the discounted return objective that is common in RL research. In this appendix, we extend these results to the finite-horizon, non-discounted setting that is common in contemporary uses of RL to train reasoning behavior in LLMs.
B.1 RL in Time-Homogenous Finite-Horizon Markov Decision Processes
In this section and the following, we use $[k]$ to denote the set of integers $\{1,...,k\}$. Formally, a time-homogenous, fixed-horizon MDP is a tuple, $\langle\mathcal{S},\mathcal{A},l,p,r\rangle$, where $\mathcal{S}$ is the environment state space, $\mathcal{A}$ is the set of actions that the agent can take to influence the environment state, $l$ is the maximum length of an episode, $p:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is a stochastic state transition function, and $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a scalar-valued reward function. When the agent is in state $s$ at timestep $t$ and takes action $a$ then it receives $r(s,a)$ and transition to a new state $S^{\prime}\sim p(\cdot|s,a)$ and $t$ is incremented by 1. The process repeats from $S^{\prime}$ until either a terminal state, $s_{\infty}$, is reached or $t$ reaches $l$. The agent selects actions according to a policy, $\pi:\mathcal{S}\times[l]\rightarrow\Delta(\mathcal{A})$. The time-dependent value of using a policy from a particular state, $s$, is defined to be $v_{\pi}(s,t)\coloneqq\mathbf{E}_{\pi}[\sum_{i=t}^{l-1}r(s_{i},a_{i})|s_{t}=s,a_{i}\sim\pi(\cdot|s_{i},i)]$. In finite-horizon RL, the agent’s objective is to find a policy that maximizes $v_{\pi}(s,t)$ in all states and for all time-steps though, in practice, we could just focuse on maximizing $v_{\pi}(s,0)$ for all possible initial states.
B.2 Time-Homogenous Finite-Horizon Thought MDPs
We formally define a finite-horizon thought MDP as the tuple $\langle{\mathcal{S}},{\mathcal{A}},l,p,r,{\mathcal{T}},{\mathcal{C}},p_{T}\rangle$, where ${\mathcal{S}},{\mathcal{A}},p,r,l$ are defined as they are for standard finite-horizon MDPs. We add $\mathcal{T}$} as the set of thought states, $\mathcal{C}$} as the set of thought actions, and $p_{T}:{\mathcal{S}}\times{\mathcal{T}}\times{\mathcal{C}}\rightarrow\Delta({\mathcal{T}})$ as the thought transition function. Recalling our informal definition of thinking from the main text, we emphasize that thought states and actions do not affect environment state transitions and rewards. However, they do affect the transition of time as we assume that the episode time-step will still be incremented when the agent takes a thought action.
We define policies for thought MDPs as in the main text except now with a time dependency: $\pi:{\mathcal{S}}\times{\mathcal{T}}\times[l]\rightarrow\Delta({\mathcal{A}}\cup{\mathcal{C}})$. In this section, we will use the notation $\pi(\tau)$ to refer to the agent’s state- and time-dependent policy with the thought state fixed at $\tau$.
In a thought MDP, interaction proceeds as follows. Episodes begin in timestep $t=0$, initial environment state, $s_{0}$, and thought state, $\tau_{0}$, and the agent chooses either an environment action or thought action according to its policy. In either case, we increment $t$ to $t=1$. If $a_{0}\in{\mathcal{A}}$ then the environment state, $s_{1}$, at the next time-step is sampled from $p(s_{0},a_{0})$, the thought state $\tau_{1}$ keeps the value of $\tau_{0}$, and the agent receives reward $r_{0}\coloneqq r(s_{0},a_{0})$. Conversely, if $a_{0}\in{\mathcal{C}}$, then $s_{1}$ keeps the value of $s_{0}$, $\tau_{1}$ is sampled from $p_{T}(s_{0},\tau_{0},a_{0})$, and the agent receives $r_{0}=0$. Then the action-selection process repeats until either the agent reaches $s_{\infty}$ or $t=l$. For thought MDPs, we define the value function to be $v_{\pi}(s,\tau,t)\coloneqq\mathbf{E}_{\pi}[\sum_{i=t}^{l-1}r(s_{i},x_{i})|s_{t}=s,\tau_{t}=\tau],$ with $r(s_{i},x_{i})=0$ if $x_{i}\in{\mathcal{C}}.$
We next adapt the key assumptions of the main paper to the finite-horizon setting. First, we assume that thought state transitions are deterministic in order to simplify our analysis and because it is also the case for LLM agents.
Assumption 9 (Deterministic Thought Transitions).
$\forall s\in{\mathcal{S}},\tau\in{\mathcal{T}},c\in{\mathcal{C}}$, $p_{{\mathcal{T}}}({\tau}^{\prime}|s,\tau,c)=1$ for one and only one ${\tau}^{\prime}\in{\mathcal{T}}$. We will write $p_{{\mathcal{T}}}(s,\tau,c)\in{\mathcal{T}}$ to denote the thought state that results from taking $c$ in $(s,\tau)$.
Assumption 10 (Non-negative Rewards).
$\forall s\in{\mathcal{S}},a\in\mathcal{A}$, $r(s,a)\geq 0$.
Modeling of Time
As in the discounted case, the fixed-horizon case raises questions about the passage of time. Just as in the main paper we assumed that discounting is applied equally between thinking and non-thinking time-steps, here we assume that both thought and environment actions take a single time-step. This assumption discourages thought actions as they cost the agent a chance to directly or indirectly obtain reward through an environment action. Potentially, we could consider varying times for thought and environment actions, however, this raises additional complexity that we will not consider further here. As before, the proposed model assumes that the environment state remains constant while the agent takes thought actions except for the passage of time. Again, for simplicity, we will focus on the static environment case, but note this issue as another interesting direction to refine the model. Finally, we only consider time-homogenous finite-horizon thought MDPs. In time-inhomogenous MDPs, the state transition probabilities and reward can differ across time-steps. In a time-inhomogenous finite-horizon thought MDP, thought actions might subtly change the state of the environment for the agent by allowing it to wait for more favorable state transition probabilities. A possible resolution would be to not advance the time-step when the agent selects a thought action. However, that resolution could break the finite-horizon assumption. For simplicity, we will only consider the time-homogenous finite-horizon setting.
Optimality
Finite-horizon thought MDPs are MDPs with state space ${\mathcal{S}}\times{\mathcal{T}}\times[l]$ and action space ${\mathcal{A}}\cup{\mathcal{C}}$ and consequently there will always be at least one deterministic optimal policy [Sutton and Barto, 2018]. Furthermore, we can show that this optimal policy will never take a thought action if we assume that environment actions are preferred to thought actions when the expected return is otherwise equal. The tie-breaking assumption means that thought actions are never strictly preferred to environment actions, though there may be situations when either type of action is equally preferred. Equal preference results in states from which it is impossible to obtain more future reward.
Proposition 11.
Assume that ties (w.r.t. expected reward-to-go) for the optimal action are broken in favor of environment actions. Then, any optimal policy, $\pi^{\star}$, for a thought MDP does not take thought actions: $\pi^{\star}(s,\tau,t)\in{\mathcal{A}}$, $\forall s\in{\mathcal{S}},\tau\in{\mathcal{T}},t\in[l]$.
Proof.
The proof is by induction. First, as the base case, consider the final time-step l1l-1. Under the optimal policy, the agent receives $r(s,\pi^{*}(s,\tau,l-1))$ in $(s,\tau)$ and then interaction terminates. Since $r(s,\pi^{*}(s,\tau,l-1))=0$ if $\pi^{*}(\tau,\tau,l-1)\in{\mathcal{C}}$ and $r(s,\pi^{*}(s,\tau,l-1))\geq 0$ if $\pi^{*}(\tau,\tau,l-1)\in{\mathcal{A}}$ then a thought action will never be strictly preferred to an environment action. Under the assumption that ties are broken in favor of environment actions, the optimal policy will select an environment action. Now, assume for arbitrary non-initial time-step $t>0$ that the optimal policy only takes environment actions for all $(s,\tau)$ and for all subsequent time-steps $t^{\prime}\geq t$. We need to show that $\pi^{*}(s,\tau,t-1)\in{\mathcal{A}}$. We will prove this step by contradiction. Suppose that $\pi^{*}(s,\tau,t-1)\in{\mathcal{C}}$. Then, after the agent takes $\pi^{*}(s,\tau,t-1)$, it transitions to $(s,{\tau}^{\prime})$ and will then only take environment actions until termination (by our inductive hypothesis). However, if the agent changed $\pi^{*}(s,\tau,t-1)$ to take $\pi^{(}s,{\tau}^{\prime},t)$ then it would get the same expected total reward over the next ltl-t time-steps and end up in some state, ${s}_{l-1}$ at time-step l1l-1 from which it would have one more chance to take an environment action and gain non-zero reward. Consequently, changing $\pi^{*}(s,\tau,t-1)$ to this action would produce at least as high a return, which implies that $\pi^{*}(s,\tau,t-1)\notin{\mathcal{C}}$ if $\pi^{*}$ is the optimal policy and ties are broken in favor of environment actions.
Appendix C Policy Initialization Determines Emergence of Thinking
Finally, we extend Theorem˜5 to the finite-horizon setting.
Theorem 12.
Let $\pi$ be a policy in a thought MDP such that $\pi(s,\tau,t)\in{\mathcal{A}}$ for some environment state $s$, thought state $\tau$, and time-step, $t$. If the policy improvement step of policy iteration sets $\pi^{\prime}(s,\tau,t)\leftarrow c$ for $c\in{\mathcal{C}}$, then $v_{\pi}(s,{\tau}^{\prime},t+1)>v_{\pi}(s,\tau,t)$, where $\tau$$^{\prime}$ is the thought state resulting from taking $c$ in $(s,\tau)$ at step $t$.
Proof.
Policy iteration sets $\pi^{\prime}(s,\tau,t)\leftarrow\arg\max_{{x}\in{\mathcal{A}}\cup{\mathcal{C}}}q_{\pi}(s,\tau,t,{x})$ where
$q_{\pi}(s,\tau,t,{x})\coloneqq\begin{cases}r(s,{x})+\mathbf{E}_{{s}^{\prime}}[v_{\pi}({s}^{\prime},\tau,t+1)]&\text{if }{x}\in{\mathcal{A}}\\ v_{\pi}(s,{\tau}^{\prime},t+1)&\text{if }{x}\in{\mathcal{C}}.\end{cases}$
$\forall a\in{\mathcal{A}},q_{\pi}(s,\tau,t,a)<v_{\pi}(s,{\tau}^{\prime},t+1).$
Since $\pi(s,\tau,t)$ was in $\mathcal{A}$} before the update we have that:
$v_{\pi}(s,\tau,t)=q_{\pi}(s,\tau,t,\pi(s,\tau,t))<v_{\pi}(s,{\tau}^{\prime},t+1).$
Thus, a thought action is selected when the policy is such that first changing the thought state to $\tau$$^{\prime}$ leads to greater expected return than taking any environment action immediately, even though changing the thought state decreases the number of remaining time-steps in which the agent can obtain reward. ∎
Appendix D Example Prompts for LLM Experiment
@@
Appendix E Extended Empirical Description of Gridworld Experiments
@@
The objective of our experiment described in Section˜6 was to create a setting where model-free RL (in this case REINFORCE) would lead to thinking as a strategy for reward maximization. A crucial piece of the set-up was to have a pre-trained model for which thought actions increase the agent’s probability of solving the task. Recall that solving the task requires the agent to navigate to the bottom right corner and then to reach the top left corner. Pre-training is designed so that the agent outputting special action ‘A’ increases the probability of it then taking the actions that lead to the bottom right and similarly so that ‘B’ increases the probability of moving to the top left. We validate the pre-training procedure by taking pre-trained models and forcing their first action to be ‘A’ and their action upon reaching the bottom right for the first time to be ‘B’. We find that the pre-training procedure leads to models where such prompting increases the probability of task success compared to simply rolling out the model. If we further prompt the model this way and also mask thought actions on every other step then the pre-trained model will solve the task on virtually every episode. This prompting and masking procedure confirmed that pre-training had primed the model for thinking to emerge as an effective strategy.
Pre-training with Misspecified Sub-Tasks
In the Gridworld domain, we pre-train the agent’s policy so as to associate the thought actions ‘A’ and ‘B’ with two sub-tasks that will be needed on the final evaluation task. Naturally, one might wonder whether thinking will still emerge if the pre-trained sub-policies are misspecified for the final learning task. To test this, we pre-train a model using the exact procedure described above but change the evaluation domain so that the complex task has each sub-goal location shifted by one cell. Though a small change, it is a sufficient change so that the strategy discovered in our main task setting will no longer completely solve the task. With this set-up, we test whether the agent can still leverage thinking to solve the task, even though triggering the pre-trained behaviors is insufficient by itself. Figure˜3 shows similar results as Figure˜2 in which the pre-trained method that uses thinking actions remains the only method to fully solve the task. Interestingly, we no longer observe any runs of the pre-trained no-think agent successfully completing the task.
(a) Success Rate
(b) Thinking Time
Figure 3: Mean learning curves for the four agents we train in the gridworld environment with misspecified sub-tasks. The vertical axis gives the success rate for first navigating to the bottom right and then the top left corner. The horizontal axis is the iteration of policy improvement (200 episodes are collected at each iteration). We run 10 trials for each learning agent and shading indicates a 95% bootstrap confidence interval. Light lines show individual training runs.